# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 16.1.1 Basic Concepts of DeFi
  • Open Access: DeFi protocols are typically open to everyone[292]. Unlike traditional financial systems where individuals may face geographical, regulatory, or identity barriers, anyone with internet access can interact with DeFi platforms.
  • Transparency: All transactions on DeFi platforms are recorded on a public blockchain[293]. This transparency ensures that the protocols are auditable and verifiable by anyone.
  • Smart Contracts: DeFi applications rely heavily on smart contracts—self-executing contracts with the terms of the agreement written directly into code[294]. These contracts automatically enforce agreements, reducing the need for human intervention and central authorities.
For instance, the Ethereum blockchain[185] is one of the most popular platforms for DeFi due to its ability to support complex smart contracts. DeFi platforms like Compound[218] and Aave[217] allow users to lend and borrow cryptocurrencies, while protocols like Uniswap[237] and SushiSwap[295] enable decentralized trading.
@@ 16.1.2 Lending Platforms and Decentralized Exchanges (DEX)
One of the key innovations in DeFi is the development of decentralized lending platforms and exchanges[296]. These systems remove the need for banks or brokerage firms by allowing users to directly lend, borrow, and trade assets in a peer-to-peer fashion.
@@
  • Compound: Compound[218] is a decentralized lending platform where users can supply or borrow assets like Ether (ETH) and stablecoins (such as DAI). The interest rates are determined algorithmically based on supply and demand.
  • Aave: Another leading lending platform, Aave[217], offers similar functionality but includes features like ”flash loans,” which allow users to borrow without collateral if the loan is repaid in the same transaction.
Decentralized Exchanges (DEX): Decentralized Exchanges (DEXs) like Uniswap[237] and SushiSwap[295] are platforms that allow users to trade cryptocurrencies directly with each other. Unlike traditional exchanges that rely on an order book to match buyers and sellers, DEXs utilize liquidity pools—collections of funds provided by users that are used for trading. The benefits of DEXs include:
@@ 16.1.3 Role and Application of Stablecoins
  • USDT (Tether)[297]: Pegged to the US dollar, Tether is one of the most widely used stablecoins in DeFi.
  • DAI[220]: DAI is a decentralized stablecoin issued by the MakerDAO protocol. It is backed by a pool of cryptocurrencies rather than fiat, making it unique in its operation.
@@ 16.2 Cross-Border Payments and Remittances
One of the most transformative applications of blockchain in fintech is its use in cross-border payments and remittances. Traditionally, international transfers involve high fees and delays due to the involvement of multiple intermediaries like banks and payment processors. Blockchain offers a faster, cheaper, and more efficient alternative[298, 299].
@@ 16.2.1 Application of Bitcoin in Cross-Border Payments
Bitcoin, the first and most well-known cryptocurrency, is increasingly being used for cross-border payments[300]. Due to its decentralized nature, Bitcoin allows individuals to send funds across borders without relying on banks or remittance services.
@@
A notable example is the use of Bitcoin in countries with unstable local currencies[301] or where access to traditional banking is limited. Individuals can use Bitcoin to send or receive money across borders, bypassing expensive and slow conventional systems.
@@ 16.2.2 Ripple and XRP as Cross-Border Payment Solutions
Ripple[302] is a blockchain-based payment system designed specifically for fast, low-cost cross-border transactions. Its native cryptocurrency, XRP[303], serves as a bridge currency for financial institutions, enabling liquidity without the need to hold pre-funded accounts in different currencies.
@@
  • Bank Partnerships: Ripple has established partnerships with several major banks and financial institutions, including Santander[304] and American Express[305], to facilitate cross-border settlements using XRP.
@@ 16.3.1 Concept of Digital Currency
The rise of digital currencies has led to the development of Central Bank Digital Currencies (CBDCs)[306], where central banks explore issuing state-backed digital money.
@@ 16.3.2 Central Bank Digital Currency Experiments in Various Countries
  • China’s Digital Yuan (e-CNY)[307]: China is one of the frontrunners in CBDC development. The digital yuan is being piloted in several cities across China and aims to complement physical currency while providing a more efficient and secure payment system.
  • European Central Bank’s Digital Euro[308]: The European Central Bank (ECB) is actively researching the potential issuance of a digital euro, which would serve as a digital version of the euro for use in everyday transactions across Europe.
  • Bahamas’ Sand Dollar[309]: The Sand Dollar, launched by the Central Bank of the Bahamas, is one of the world’s first fully operational CBDCs. It aims to improve financial inclusion and streamline payments in the country.
@@ 17.1.1 Transparency and Traceability in Supply Chains
In modern supply chains, which often span multiple countries and involve various stakeholders, transparency and traceability are major challenges[310]. As products move from raw materials through different stages of production and logistics, it becomes increasingly difficult to maintain visibility into each stage of the process. Lack of transparency can result in inefficiencies, increased costs, and potential fraud[311].
@@
In addition to inefficiencies, lack of traceability can lead to fraud[312]. Unscrupulous actors may replace genuine goods with counterfeit ones or misrepresent the origin of products. This is a significant problem in industries such as pharmaceuticals, where counterfeit drugs can pose serious health risks, or luxury goods, where counterfeit products can harm brand reputation.
@@ 17.1.2 Trust Issues in Supply Chain Management
Trust is another significant issue in supply chain management. In a globalized economy, supply chains are often comprised of multiple entities—manufacturers, suppliers, logistics providers, and retailers—many of whom do not fully trust each other. This lack of trust can lead to delays, errors, and disputes[313].
@@
Traditionally, intermediaries such as banks or third-party verification agencies are required to build trust between these parties[314]. These intermediaries validate transactions, confirm deliveries, and manage payments. However, this reliance on intermediaries adds cost and complexity to the process. Moreover, intermediaries themselves can be subject to fraud or corruption, introducing new vulnerabilities into the system.
@@ 17.2.1 Role of Blockchain in Goods Tracking
Blockchain technology can significantly enhance the tracking of goods in supply chains by providing a transparent, decentralized, and tamper-proof system. With blockchain, every transaction, movement, or change of ownership of goods can be recorded in a digital ledger that is accessible to all stakeholders[315].
@@ 17.2.2 Immutability of Supply Chain Data
One of the most important features of blockchain is its immutability. Once data is recorded on a blockchain, it cannot be altered or deleted. This feature is particularly valuable in supply chain management, where the integrity of data is critical for ensuring transparency, trust, and accountability[316].
@@
Immutability also helps in regulatory compliance. For industries such as pharmaceuticals or food, where traceability is legally required, blockchain provides an unchangeable record of the movement of goods[317]. This not only simplifies audits but also helps companies prove compliance with regulations, reducing the risk of fines or legal action.
@@ 17.3.1 Using Blockchain to Ensure Product Authenticity
Counterfeiting is a major issue in several industries, including luxury goods, electronics, and pharmaceuticals. Blockchain can be used to verify the authenticity of products by providing a permanent, verifiable record of their origin and history[318].
@@
In the pharmaceutical industry, blockchain can be used to track the movement of drugs from the manufacturer to the pharmacy[319]. This helps prevent counterfeit drugs from entering the supply chain, as each step of the drug’s journey is recorded on the blockchain and can be verified by healthcare providers and consumers.
@@ 17.3.2 End-to-End Traceability from Raw Materials to Consumers
Blockchain enables end-to-end traceability of products, from the sourcing of raw materials to the final delivery to the consumer[320]. This level of transparency helps build consumer trust and ensures product quality.
@@
Consumers can access this information through a mobile app or website by scanning a QR code on the product packaging. This allows them to verify the origin of the product and check for any certifications, such as organic or fair trade. In cases of foodborne illness outbreaks, blockchain can help trace the contaminated product back to its source more quickly, reducing the impact of recalls[321].
@@ 18.1.1 Concept and Principles of DID
Decentralized Identity (DID)[322] is a groundbreaking concept that allows individuals to have full control over their digital identities without needing to rely on centralized authorities, such as governments or large corporations. Traditional identity systems often depend on central entities, which can lead to security risks, privacy issues, and a lack of control for individuals. DID, on the other hand, ensures that each person owns and manages their identity autonomously.
@@ 18.1.2 Role of DID in Data Privacy Protection
One of the key benefits of DID systems is the enhanced protection of user privacy[323]. In traditional systems, users often have to share personal data (such as email, address, or full identification details) with multiple service providers, each of whom stores that information centrally. This increases the risk of data breaches, identity theft, and unauthorized access.
In a DID system, individuals can control exactly what information they share with service providers. For instance, if a service requires proof of age, a user could share a cryptographically signed proof that they are over a certain age, without needing to share their full birthdate or identity. This concept is known as ”selective disclosure.”[324]
@@ 18.2.2 Ensuring the Security of Digital Identity
    Encryption: Cryptographic encryption techniques protect data associated with DIDs. Public-key cryptography is used to ensure that only the owner of a DID can control it[325]. The private key is held by the user, and the public key is recorded on the blockchain.
@@
    Cryptographic Keys: Each DID is associated with a pair of cryptographic keys (public and private keys). The private key remains with the user, and the public key is shared with others for verification[326]. For example, when a user logs in to a service using their DID, the service can verify the user’s identity by checking their public key on the blockchain.
@@
    Decentralized Verification: Unlike traditional systems where a single entity verifies identity, blockchain allows decentralized verification[327]. Multiple nodes in the blockchain network can validate the authenticity of an identity or credential, reducing the risk of fraud or unauthorized changes.
@@ 19.1.1 Indivisibility and Uniqueness of NFTs
Non-Fungible Tokens, or NFTs, are a type of digital asset that represent ownership of unique items on the blockchain. Two key characteristics of NFTs are their indivisibility and uniqueness[328]. Unlike cryptocurrencies such as Bitcoin or Ethereum, which can be divided into smaller units and are fungible (interchangeable), NFTs are unique and cannot be split into smaller parts[329].
@@ 19.1.2 Differences Between NFTs and ERC-20 Tokens
NFTs are often compared to ERC-20 tokens[330], another popular type of token used on the Ethereum blockchain. However, there are significant differences between the two:
  • Fungibility: ERC-20 tokens are fungible[331], meaning one token is identical to another and can be exchanged without any difference. For instance, one DAI token is equivalent to another DAI token in every way[332]. On the other hand, NFTs are non-fungible, meaning each token is unique and not interchangeable with another NFT.
  • Use Cases: ERC-20 tokens are commonly used for digital currencies, governance tokens, and other assets where interchangeability is required[331]. NFTs, by contrast, are used to represent unique assets such as digital art, collectibles, virtual real estate, or music[333].
  • Market Dynamics: The market for ERC-20 tokens behaves like traditional financial markets where tokens can be traded based on supply and demand with uniform pricing[334]. NFTs, however, have markets more akin to the art world, where the value of each asset is determined individually based on rarity, creator reputation, and demand[335].
@@ 19.2.1 ERC-721 Standard
The ERC-721 standard defines the technical foundation for NFTs on the Ethereum blockchain[223]. This standard outlines the methods that must be implemented for a contract to handle NFTs. The most critical feature of the ERC-721 standard is its ability to ensure the interoperability and uniqueness of NFTs across different platforms and applications.
@@ 19.2.2 NFT Creation and Transaction Process
The process of creating (or minting) an NFT involves deploying a smart contract that adheres to the ERC-721 standard. The metadata associated with the NFT, such as its name, description, and link to the digital file (e.g., an image or video), is also stored in the blockchain or in decentralized storage systems like IPFS[336].
@@ 19.3.1 Digital Art
For example, consider the digital artist Beeple[337], who sold his NFT artwork ”Everydays: The First 5000 Days” for over \$69 million[338]. This sale exemplifies the growing interest in digital art NFTs and how they can provide new revenue streams for artists.

### v3

@@ 16.1.1 Basic Concepts of DeFi
  • Open Access: DeFi protocols are typically open to everyone[Salami2021ChallengesAA]. Unlike traditional financial systems where individuals may face geographical, regulatory, or identity barriers, anyone with internet access can interact with DeFi platforms.
  • Transparency: All transactions on DeFi platforms are recorded on a public blockchain[secSECgovStatement]. This transparency ensures that the protocols are auditable and verifiable by anyone.
  • Smart Contracts: DeFi applications rely heavily on smart contracts—self-executing contracts with the terms of the agreement written directly into code[John2023SmartCA]. These contracts automatically enforce agreements, reducing the need for human intervention and central authorities.
For instance, the Ethereum blockchain[ethereumCompleteGuide] is one of the most popular platforms for DeFi due to its ability to support complex smart contracts. DeFi platforms like Compound[compoundv2CompoundUnderstanding] and Aave[aaveAave] allow users to lend and borrow cryptocurrencies, while protocols like Uniswap[uniswapUniswapInterface] and SushiSwap[sushisvapSushiSwapAdvanced] enable decentralized trading.
@@ 16.1.2 Lending Platforms and Decentralized Exchanges (DEX)
One of the key innovations in DeFi is the development of decentralized lending platforms and exchanges[Carapella2022DecentralizedF]. These systems remove the need for banks or brokerage firms by allowing users to directly lend, borrow, and trade assets in a peer-to-peer fashion.
@@
  • Compound: Compound[compoundv2CompoundUnderstanding] is a decentralized lending platform where users can supply or borrow assets like Ether (ETH) and stablecoins (such as DAI). The interest rates are determined algorithmically based on supply and demand.
  • Aave: Another leading lending platform, Aave[aaveAave], offers similar functionality but includes features like ”flash loans,” which allow users to borrow without collateral if the loan is repaid in the same transaction.
Decentralized Exchanges (DEX): Decentralized Exchanges (DEXs) like Uniswap[uniswapUniswapInterface] and SushiSwap[sushisvapSushiSwapAdvanced] are platforms that allow users to trade cryptocurrencies directly with each other. Unlike traditional exchanges that rely on an order book to match buyers and sellers, DEXs utilize liquidity pools—collections of funds provided by users that are used for trading. The benefits of DEXs include:
@@ 16.1.3 Role and Application of Stablecoins
  • USDT (Tether)[tether]: Pegged to the US dollar, Tether is one of the most widely used stablecoins in DeFi.
  • DAI[makerdaoMakerDAOUnbiased]: DAI is a decentralized stablecoin issued by the MakerDAO protocol. It is backed by a pool of cryptocurrencies rather than fiat, making it unique in its operation.
@@ 16.2 Cross-Border Payments and Remittances
One of the most transformative applications of blockchain in fintech is its use in cross-border payments and remittances. Traditionally, international transfers involve high fees and delays due to the involvement of multiple intermediaries like banks and payment processors. Blockchain offers a faster, cheaper, and more efficient alternative[forbesCouncilPost, Qiu2018RippleVS].
@@ 16.2.1 Application of Bitcoin in Cross-Border Payments
Bitcoin, the first and most well-known cryptocurrency, is increasingly being used for cross-border payments[bitcoinBitcoinOpen]. Due to its decentralized nature, Bitcoin allows individuals to send funds across borders without relying on banks or remittance services.
@@
A notable example is the use of Bitcoin in countries with unstable local currencies[he2024cryptocurrency] or where access to traditional banking is limited. Individuals can use Bitcoin to send or receive money across borders, bypassing expensive and slow conventional systems.
@@ 16.2.2 Ripple and XRP as Cross-Border Payment Solutions
Ripple[rippleGlobalPayments] is a blockchain-based payment system designed specifically for fast, low-cost cross-border transactions. Its native cryptocurrency, XRP[rippleDigitalAsset], serves as a bridge currency for financial institutions, enabling liquidity without the need to hold pre-funded accounts in different currencies.
@@
  • Bank Partnerships: Ripple has established partnerships with several major banks and financial institutions, including Santander[rippleSantanderPartners] and American Express[rippleAmericanExpress], to facilitate cross-border settlements using XRP.
@@ 16.3.1 Concept of Digital Currency
The rise of digital currencies has led to the development of Central Bank Digital Currencies (CBDCs)[Griffoli2018CastingLO], where central banks explore issuing state-backed digital money.
@@ 16.3.2 Central Bank Digital Currency Experiments in Various Countries
  • China’s Digital Yuan (e-CNY)[cnbcChinaGiven]: China is one of the frontrunners in CBDC development. The digital yuan is being piloted in several cities across China and aims to complement physical currency while providing a more efficient and secure payment system.
  • European Central Bank’s Digital Euro[europaDigitalEuro]: The European Central Bank (ECB) is actively researching the potential issuance of a digital euro, which would serve as a digital version of the euro for use in everyday transactions across Europe.
  • Bahamas’ Sand Dollar[sanddollar]: The Sand Dollar, launched by the Central Bank of the Bahamas, is one of the world’s first fully operational CBDCs. It aims to improve financial inclusion and streamline payments in the country.
@@ 17.1.1 Transparency and Traceability in Supply Chains
In modern supply chains, which often span multiple countries and involve various stakeholders, transparency and traceability are major challenges[Saxena2023BlockchainFS]. As products move from raw materials through different stages of production and logistics, it becomes increasingly difficult to maintain visibility into each stage of the process. Lack of transparency can result in inefficiencies, increased costs, and potential fraud[Patterson2018UnderstandingAM].
@@
In addition to inefficiencies, lack of traceability can lead to fraud[DuHadway2022LightID]. Unscrupulous actors may replace genuine goods with counterfeit ones or misrepresent the origin of products. This is a significant problem in industries such as pharmaceuticals, where counterfeit drugs can pose serious health risks, or luxury goods, where counterfeit products can harm brand reputation.
@@ 17.1.2 Trust Issues in Supply Chain Management
Trust is another significant issue in supply chain management. In a globalized economy, supply chains are often comprised of multiple entities—manufacturers, suppliers, logistics providers, and retailers—many of whom do not fully trust each other. This lack of trust can lead to delays, errors, and disputes[Imeri2019BlockchainAO].
@@
Traditionally, intermediaries such as banks or third-party verification agencies are required to build trust between these parties[Yaksick2019OvercomingSC]. These intermediaries validate transactions, confirm deliveries, and manage payments. However, this reliance on intermediaries adds cost and complexity to the process. Moreover, intermediaries themselves can be subject to fraud or corruption, introducing new vulnerabilities into the system.
@@ 17.2.1 Role of Blockchain in Goods Tracking
Blockchain technology can significantly enhance the tracking of goods in supply chains by providing a transparent, decentralized, and tamper-proof system. With blockchain, every transaction, movement, or change of ownership of goods can be recorded in a digital ledger that is accessible to all stakeholders[Reddy2019EnhancingSC].
@@ 17.2.2 Immutability of Supply Chain Data
One of the most important features of blockchain is its immutability. Once data is recorded on a blockchain, it cannot be altered or deleted. This feature is particularly valuable in supply chain management, where the integrity of data is critical for ensuring transparency, trust, and accountability[Wu2019DataMI].
@@
Immutability also helps in regulatory compliance. For industries such as pharmaceuticals or food, where traceability is legally required, blockchain provides an unchangeable record of the movement of goods[Scott2018EvaluatingFO]. This not only simplifies audits but also helps companies prove compliance with regulations, reducing the risk of fines or legal action.
@@ 17.3.1 Using Blockchain to Ensure Product Authenticity
Counterfeiting is a major issue in several industries, including luxury goods, electronics, and pharmaceuticals. Blockchain can be used to verify the authenticity of products by providing a permanent, verifiable record of their origin and history[Thakkar2021ApplicationFC].
@@
In the pharmaceutical industry, blockchain can be used to track the movement of drugs from the manufacturer to the pharmacy[Haq2018BlockchainTI]. This helps prevent counterfeit drugs from entering the supply chain, as each step of the drug’s journey is recorded on the blockchain and can be verified by healthcare providers and consumers.
@@ 17.3.2 End-to-End Traceability from Raw Materials to Consumers
Blockchain enables end-to-end traceability of products, from the sourcing of raw materials to the final delivery to the consumer[Cruz2020BlockchainbasedTP]. This level of transparency helps build consumer trust and ensures product quality.
@@
Consumers can access this information through a mobile app or website by scanning a QR code on the product packaging. This allows them to verify the origin of the product and check for any certifications, such as organic or fair trade. In cases of foodborne illness outbreaks, blockchain can help trace the contaminated product back to its source more quickly, reducing the impact of recalls[Oriekhoe2024ENSURINGGF].
@@ 18.1.1 Concept and Principles of DID
Decentralized Identity (DID)[NassarKyriakidou2023DecentralizedIW] is a groundbreaking concept that allows individuals to have full control over their digital identities without needing to rely on centralized authorities, such as governments or large corporations. Traditional identity systems often depend on central entities, which can lead to security risks, privacy issues, and a lack of control for individuals. DID, on the other hand, ensures that each person owns and manages their identity autonomously.
@@ 18.1.2 Role of DID in Data Privacy Protection
One of the key benefits of DID systems is the enhanced protection of user privacy[2019EIDASSS]. In traditional systems, users often have to share personal data (such as email, address, or full identification details) with multiple service providers, each of whom stores that information centrally. This increases the risk of data breaches, identity theft, and unauthorized access.
In a DID system, individuals can control exactly what information they share with service providers. For instance, if a service requires proof of age, a user could share a cryptographically signed proof that they are over a certain age, without needing to share their full birthdate or identity. This concept is known as ”selective disclosure.”[Kim2021AnalysisOT]
@@ 18.2.2 Ensuring the Security of Digital Identity
    Encryption: Cryptographic encryption techniques protect data associated with DIDs. Public-key cryptography is used to ensure that only the owner of a DID can control it[Min2021ASO]. The private key is held by the user, and the public key is recorded on the blockchain.
@@
    Cryptographic Keys: Each DID is associated with a pair of cryptographic keys (public and private keys). The private key remains with the user, and the public key is shared with others for verification[Park2022ANA]. For example, when a user logs in to a service using their DID, the service can verify the user’s identity by checking their public key on the blockchain.
@@
    Decentralized Verification: Unlike traditional systems where a single entity verifies identity, blockchain allows decentralized verification[Thorve2022DecentralizedIM]. Multiple nodes in the blockchain network can validate the authenticity of an identity or credential, reducing the risk of fraud or unauthorized changes.
@@ 19.1.1 Indivisibility and Uniqueness of NFTs
Non-Fungible Tokens, or NFTs, are a type of digital asset that represent ownership of unique items on the blockchain. Two key characteristics of NFTs are their indivisibility and uniqueness[Idelberger2022NonfungibleT]. Unlike cryptocurrencies such as Bitcoin or Ethereum, which can be divided into smaller units and are fungible (interchangeable), NFTs are unique and cannot be split into smaller parts[Martin2021ATD].
@@ 19.1.2 Differences Between NFTs and ERC-20 Tokens
NFTs are often compared to ERC-20 tokens[Parham2022NonfungibleTP], another popular type of token used on the Ethereum blockchain. However, there are significant differences between the two:
  • Fungibility: ERC-20 tokens are fungible[Zhang2023TheAO], meaning one token is identical to another and can be exchanged without any difference. For instance, one DAI token is equivalent to another DAI token in every way[Feng2024ModelingAA]. On the other hand, NFTs are non-fungible, meaning each token is unique and not interchangeable with another NFT.
  • Use Cases: ERC-20 tokens are commonly used for digital currencies, governance tokens, and other assets where interchangeability is required[Zhang2023TheAO]. NFTs, by contrast, are used to represent unique assets such as digital art, collectibles, virtual real estate, or music[Gupta2023ReviewingTR].
  • Market Dynamics: The market for ERC-20 tokens behaves like traditional financial markets where tokens can be traded based on supply and demand with uniform pricing[Somin2018NetworkAO]. NFTs, however, have markets more akin to the art world, where the value of each asset is determined individually based on rarity, creator reputation, and demand[Mekacher2022HowRS].
@@ 19.2.1 ERC-721 Standard
The ERC-721 standard defines the technical foundation for NFTs on the Ethereum blockchain[ethereumERC721NonFungible]. This standard outlines the methods that must be implemented for a contract to handle NFTs. The most critical feature of the ERC-721 standard is its ability to ensure the interoperability and uniqueness of NFTs across different platforms and applications.
@@ 19.2.2 NFT Creation and Transaction Process
The process of creating (or minting) an NFT involves deploying a smart contract that adheres to the ERC-721 standard. The metadata associated with the NFT, such as its name, description, and link to the digital file (e.g., an image or video), is also stored in the blockchain or in decentralized storage systems like IPFS[ipfsOpenSystem].
@@ 19.3.1 Digital Art
For example, consider the digital artist Beeple[beeplecrapHOMEBEEPLE], who sold his NFT artwork ”Everydays: The First 5000 Days” for over \$69 million[christiesBeeple1981]. This sale exemplifies the growing interest in digital art NFTs and how they can provide new revenue streams for artists.